OpenStudy (anonymous):

Please help! Simplify complex fraction: 29+(2/3)x/72+x

OpenStudy (anonymous):

Complex as in "difficult" or complex as in "has an imaginary part"?

OpenStudy (anonymous):

Complex Fraction: Fraction that has a fraction as part of its numerator, denominator, or both. In other words, difficult. Can you help me?

OpenStudy (anonymous):

Sure thing. \[\frac{29+\frac{2}{3}x}{72+x}\] Get a common denominator for the top part. \[29+\frac{2}{3}x=\frac{29*3}{3}+\frac{2x}{3}=\frac{87}{3}+\frac{2x}{3}=\frac{2x+87}{3}\] Our fraction is now\[\frac{\frac{2x+87}{3}}{72+x}=\frac{2x+87}{3}\frac{1}{x+72}\] \[=\frac{2x+87}{3x+216}\] Might want to have someone check my work. I may have made a mistake in there.

OpenStudy (anonymous):

Thank you a bunch! :)

OpenStudy (anonymous):

No problem :)